OpenStudy (matt6288):

Two positive numbers have a difference of 8 and a product of 33. What are these numbers?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x - y = 8 }\) \(\Large \color{purple}{\rightarrow xy = 33 }\) Solve the system of equations.

OpenStudy (matt6288):

# A. 4 and 12 # B. 3 and 12 # C. 4 and 11 # D. 3 and 11

OpenStudy (amistre64):

2x - 2y = 8 4xy = 33 might be a little better

OpenStudy (amistre64):

for some reason i read that as 2 even numbers :/ i need new glasses

Parth (parthkohli):

You might even check it from the choices.

Parth (parthkohli):

12 - 4 = 8 12 * 4 = 48 ^ NO 12 - 3 = 9 ^ NO 11 - 4 = 7 ^ NO 11 - 3 = 8 11 * 3 = 33 ^ YES!!!